# Evaluating language models as risk scores

### v1

@@ Acknowledgments
This appendix section shows additional results and corresponding plots to support the insights presented in Section 4. Section A.1 shows results on four extra benchmark tasks made available with folktexts. Section A.2 extends the discussion on subgroup calibration and algorithmic fairness on the ACSIncome task. Section A.3 goes further in-depth on how to use folktexts to control data uncertainty in the benchmark prediction tasks. Finally, Section A.4 presents and discusses results on feature importance for LLM predictions.
A.1 Results on additional benchmark tasks
@@
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.17 0.19 0.85 0.73 0.78 0.29
Llama 3 70B 0.25 0.26 0.82 0.51 0.74 0.12
Llama 3 8B (it) 0.07 0.19 0.79 0.74 0.74 0.22
Llama 3 8B 0.34 0.34 0.75 0.45 0.68 0.06
Mixtral 8x22B (it) 0.38 0.39 0.60 0.51 0.60 0.26
Mixtral 8x22B 0.21 0.24 0.86 0.52 0.77 0.08
Mixtral 8x7B (it) 0.22 0.23 0.83 0.73 0.75 0.44
Mixtral 8x7B 0.29 0.31 0.81 0.45 0.74 0.06
Mistral 7B (it) 0.35 0.36 0.72 0.63 0.66 0.49
Mistral 7B 0.26 0.30 0.76 0.45 0.69 0.04
Yi 34B (chat) 0.14 0.21 0.79 0.69 0.73 0.20
Yi 34B 0.08 0.23 0.70 0.62 0.63 0.11
Gemma 7B (it) 0.36 0.38 0.59 0.58 0.59 0.44
Gemma 7B 0.16 0.25 0.65 0.48 0.64 0.13
Gemma 2B (it) 0.38 0.41 0.42 0.46 0.48 0.23
Gemma 2B 0.01 0.24 0.63 0.54 0.61 0.03
LR 0.02 0.15 0.86 0.78 - -
XGBoost 0.00 0.12 0.91 0.83 - -
@@
Table A1: Zero-shot LLM results on the ACSEmployment benchmark task, together with supervised learning baselines fitted on 2.9M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.20 0.25 0.57 0.61 0.48 0.18
Llama 3 70B 0.22 0.24 0.55 0.67 0.53 0.04
Llama 3 8B (it) 0.15 0.22 0.56 0.70 0.53 0.17
Llama 3 8B 0.10 0.20 0.55 0.73 0.63 0.04
Mixtral 8x22B (it) 0.41 0.40 0.51 0.40 0.50 0.24
Mixtral 8x22B 0.11 0.20 0.55 0.73 0.63 0.02
Mixtral 8x7B (it) 0.26 0.26 0.58 0.73 0.60 0.02
Mixtral 8x7B 0.14 0.21 0.57 0.73 0.52 0.03
Mistral 7B (it) 0.26 0.26 0.57 0.73 0.54 0.02
Mistral 7B 0.20 0.23 0.53 0.73 0.33 0.01
Yi 34B (chat) 0.09 0.20 0.56 0.72 0.59 0.15
Yi 34B 0.07 0.20 0.57 0.73 0.63 0.08
Gemma 7B (it) 0.25 0.26 0.58 0.73 0.67 0.11
Gemma 7B 0.41 0.37 0.50 0.39 0.41 0.10
Gemma 2B (it) 0.73 0.73 0.52 0.27 0.40 0.00
Gemma 2B 0.25 0.26 0.51 0.39 0.45 0.02
LR 0.02 0.19 0.61 0.74 - -
XGBoost 0.00 0.16 0.74 0.76 - -
@@
Table A2: Zero-shot LLM results on the ACSMobility benchmark task, together with supervised learning baselines fitted on 0.6M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.15 0.24 0.70 0.60 0.61 0.12
Llama 3 70B 0.09 0.24 0.67 0.55 0.61 0.05
Llama 3 8B (it) 0.19 0.28 0.60 0.57 0.57 0.11
Llama 3 8B 0.08 0.25 0.53 0.56 0.54 0.03
Mixtral 8x22B (it) 0.32 0.33 0.66 0.59 0.61 0.15
Mixtral 8x22B 0.20 0.28 0.63 0.44 0.54 0.04
Mixtral 8x7B (it) 0.45 0.45 0.66 0.52 0.55 0.29
Mixtral 8x7B 0.28 0.32 0.60 0.44 0.56 0.03
Mistral 7B (it) 0.41 0.42 0.59 0.57 0.53 0.19
Mistral 7B 0.05 0.25 0.57 0.56 0.55 0.02
Yi 34B (chat) 0.35 0.36 0.65 0.56 0.52 0.02
Yi 34B 0.08 0.24 0.62 0.56 0.52 0.06
Gemma 7B (it) 0.42 0.43 0.53 0.56 0.50 0.02
Gemma 7B 0.04 0.24 0.61 0.58 0.57 0.02
Gemma 2B (it) 0.34 0.36 0.49 0.56 0.54 0.03
Gemma 2B 0.09 0.26 0.48 0.44 0.45 0.01
LR 0.04 0.24 0.58 0.56 - -
XGBoost 0.02 0.19 0.77 0.70 - -
@@
Table A3: Zero-shot LLM results on the ACSTravelTime benchmark task, together with supervised learning baselines fitted on 1.3M samples. See caption of Table 1 for more details.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.16 0.21 0.69 0.75 0.49 0.20
Llama 3 70B 0.18 0.22 0.67 0.63 0.50 0.07
Llama 3 8B (it) 0.11 0.21 0.59 0.71 0.71 0.19
Llama 3 8B 0.41 0.38 0.55 0.30 0.50 0.02
Mixtral 8x22B (it) 0.24 0.25 0.70 0.72 0.67 0.17
Mixtral 8x22B 0.32 0.30 0.59 0.30 0.56 0.05
Mixtral 8x7B (it) 0.20 0.23 0.70 0.74 0.72 0.35
Mixtral 8x7B 0.41 0.37 0.57 0.30 0.38 0.03
Mistral 7B (it) 0.30 0.30 0.61 0.70 0.68 0.01
Mistral 7B 0.29 0.30 0.45 0.30 0.70 0.02
Yi 34B (chat) 0.06 0.19 0.67 0.74 0.68 0.13
Yi 34B 0.04 0.21 0.59 0.70 0.56 0.07
Gemma 7B (it) 0.30 0.34 0.46 0.50 0.32 0.32
Gemma 7B 0.15 0.23 0.49 0.49 0.37 0.10
Gemma 2B (it) 0.70 0.70 0.54 0.30 0.33 0.00
Gemma 2B 0.26 0.28 0.54 0.30 0.62 0.01
LR 0.03 0.19 0.70 0.73 - -
XGBoost 0.00 0.14 0.84 0.80 - -
@@
Table A4: Zero-shot LLM results on the ACSPublicCoverage benchmark task, together with supervised learning baselines fitted on 1.0M samples. See caption of Table 1 for more details.
A.2 Additional subgroup calibration results
@@
This section contains additional subgroup calibration plots on the ACSIncome task. Figure 6 shows the 2 worst offenders: Yi 34B and Mixtral 8x22B. Figures A1 and A2 show the remaining models. The same trend is visible but to a lesser extent: Instruction-tuned models show a bias towards lower scores for Black individuals, even after controlling for label prevalence.
@@
Figure A3 shows the difference between the signed calibration error (SCE) on different group pairings, on the ACSIncome task. Positive predictions ($Y=1$) correspond to the advantaged high-income class. Negative differences, $\Delta_{SCE}<0$, indicates advantaged scores for Black individuals, while positive differences, $\Delta_{SCE}>0$, indicate disadvantaged scores for Black individuals. Interestingly, while subgroup calibration curves appear similar for base models and disadvantage Black individuals for instruction-tuned models (see Figures 6A1, and A2), this is not entirely reflected on the SCE metric. Indeed, a clear-cut split is visible: base models benefit the score of Black individuals, and instruct models disadvantage the score of Black individuals. This finding could be partially explained by the fact that base models produce score distributions with low variance, and instruct models produce high-variance polarized outcomes. Specifically, two conclusions can be drawn from the score distributions produced by base models: (1) models under-estimate the score of high-income earners (which are disproportionately Asian), and (2) models over-estimate the score of low-income earners (which are disproportionately Black). This could lead base models to over-estimate the earnings of the Black population disproportionately to other groups. The opposite is true for instruction-tuned models: high-income earners see their score over-estimated, which benefits groups with a higher prevalence of high earners.
@@
Figure A1: Calibration curves for the Mistral family of models, across different race sub-populations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models.
Figure A2: Calibration curves for the Gemma family of models, across different race sub-populations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models. Note that these are the smallest models and generally perform the worst on all tasks. The plotted calibration curves are in agreement with the lackluster performance, as they show extreme score distributions that are unlike those of other LLMs. In fact, the instruction-tuned Gemma models consistently output only positive predictions, hindering their analysis.
Figure A3: Racial group bias in risk score calibration error. When comparing score bias of group $A$ with score bias of group $B$, $\Delta_{\mathrm{SCE}}=\mathrm{SCE}_{A}-\mathrm{SCE}_{B}$, positive values indicate an undue score advantage of group $A$, and negative values an undue score advantage of group $B$. Left: Difference between White and Black groups, $\mathrm{SCE}_{White}-\mathrm{SCE}_{Black}$. Right: Difference between Asian and Black groups, $\mathrm{SCE}_{Asian}-\mathrm{SCE}_{Black}$. Note that the Gemma models were omitted, as their instruct versions degenerate into strictly predicting only positive outcomes (see Figure A2). Consequently, the two instruct Gemma models are the only exceptions to the trend shown in these plots.
A.3 Varying uncertainty

### v3

@@ Acknowledgments
This appendix section shows additional results and corresponding plots to support the insights presented in Section 4. Section A.1 shows results using a chat-style verbalized numeric prompting scheme. Section A.2 shows results on four extra benchmark tasks made available with folktexts. Section A.3 extends the discussion on subgroup calibration and algorithmic fairness on the ACSIncome task. Section A.4 goes further in-depth on how to use folktexts to control data uncertainty in the benchmark prediction tasks. Finally, Section A.5 presents and discusses results on feature importance for LLM predictions.
A.1 Additional results using verbalized numeric prompting
Figure A1 shows the change in calibration error (ECE) between using multiple-choice prompting and verbalized numeric prompting, on all five benchmark tasks. Instruction-tuned models (top rows) show ECE improvements on an overwhelming majority of model/task pairs, while base models (bottom rows) show less consistent results. However, using numeric prompting comes at a consistent cost of diminished predictive power (AUC) of the risk scores, shown in Figure A2. A majority of model/task pairs have worse AUC with numeric prompting, with the exception of the employment prediction task. One potential explanation for this generalized decrease in AUC lies in the fact that numeric prompting generates a large number of tied risk scores. Figure A3 shows the score distribution of all models using numeric prompting (compare with multiple-choice prompting shown in Figure 5). While multiple-choice prompting produces a smoother continuous score distribution, numeric prompting generally results in a small set of possible uncertainty estimates. This arguably makes intuitive sense, as numeric prompting produces uncertainty estimates in discrete token space, while multiple-choice prompting produces uncertainty estimates in the continuous token-probability space.
A.2 Results on additional benchmark tasks
@@
Figure A1: Change in calibration error (ECE) when using numeric risk prompting ($\bullet$) versus multiple-choice prompting ($\blacksquare$). Instruction-tuned models (top rows) show substantial calibration improvements, while base models (bottom rows) show mixed results. Green/red arrows signal ECE improvement/degradation.
Figure A2: Change in predictive power (AUC) when using numeric risk prompting ($\bullet$) versus multiple-choice prompting ($\blacksquare$). Both instruction-tuned models (top rows) and base models (bottom rows) generally achieve worse AUC with numeric prompting. Green/red arrows signal AUC improvement/degradation, respectively.
Figure A3: Distribution of risk scores produced using numeric prompting on the ACSIncome benchmark task. A baseline score distribution that achieves $0.00$ calibration error is shown in green (XGBoost model). For each model, risk scores produced in this manner fall into only a few different possible values, contrasting with the neatly continuous distribution produced by multiple-choice prompting. This fact leads to numerous more ties among predicted risk scores, which can explain the reduced AUC performance with this prompting scheme. Nonetheless, calibration error is considerably smaller for instruction-tuned models.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.28 0.29 0.79 0.65 0.23 0.23 0.80 0.73
Mixtral 8x22B (it) 0.38 0.39 0.60 0.51 0.06 0.14 0.87 0.79
Mixtral 8x22B 0.21 0.24 0.86 0.52 0.15 0.18 0.82 0.80
Llama 3 70B (it) 0.17 0.19 0.85 0.73 0.05 0.14 0.88 0.81
Llama 3 70B 0.25 0.26 0.82 0.52 0.05 0.15 0.86 0.78
Mixtral 8x7B (it) 0.22 0.24 0.82 0.73 0.07 0.15 0.87 0.78
Mixtral 8x7B 0.30 0.31 0.81 0.45 0.08 0.17 0.81 0.73
Yi 34B (it) 0.14 0.21 0.79 0.69 0.15 0.21 0.81 0.51
Yi 34B 0.08 0.23 0.70 0.62 0.13 0.23 0.66 0.50
Llama 3 8B (it) 0.07 0.19 0.79 0.74 0.08 0.17 0.82 0.77
Llama 3 8B 0.34 0.34 0.76 0.45 0.15 0.23 0.75 0.46
Mistral 7B (it) 0.35 0.36 0.72 0.63 0.04 0.19 0.79 0.69
Mistral 7B 0.26 0.30 0.76 0.45 0.14 0.19 0.80 0.79
Gemma 7B (it) 0.36 0.38 0.59 0.58 0.04 0.22 0.71 0.60
Gemma 7B 0.15 0.25 0.65 0.48 0.35 0.38 0.50 0.51
Gemma 2B (it) 0.38 0.41 0.42 0.46 0.12 0.27 0.46 0.46
Gemma 2B 0.01 0.24 0.63 0.54 0.01 0.23 0.57 0.53
LR 0.02 0.15 0.86 0.78 0.02 0.15 0.86 0.78
XGBoost 0.00 0.12 0.91 0.83 0.00 0.12 0.91 0.83
@@
Table A1: Zero-shot LLM results on the ACSEmployment benchmark task, together with supervised learning baselines fitted on 2.9M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.26 0.26 0.57 0.73 0.22 0.25 0.49 0.73
Mixtral 8x22B (it) 0.40 0.40 0.51 0.39 0.05 0.20 0.54 0.73
Mixtral 8x22B 0.11 0.21 0.55 0.73 0.13 0.22 0.49 0.73
Llama 3 70B (it) 0.20 0.25 0.57 0.58 0.05 0.20 0.52 0.73
Llama 3 70B 0.22 0.24 0.55 0.53 0.06 0.20 0.53 0.73
Mixtral 8x7B (it) 0.26 0.26 0.58 0.73 0.11 0.21 0.51 0.73
Mixtral 8x7B 0.14 0.21 0.57 0.73 0.24 0.25 0.48 0.73
Yi 34B (it) 0.09 0.20 0.56 0.72 0.23 0.25 0.50 0.27
Yi 34B 0.07 0.20 0.57 0.73 0.15 0.23 0.52 0.44
Llama 3 8B (it) 0.15 0.22 0.56 0.70 0.11 0.21 0.49 0.73
Llama 3 8B 0.10 0.20 0.55 0.73 0.14 0.21 0.51 0.72
Mistral 7B (it) 0.26 0.26 0.57 0.73 0.17 0.23 0.49 0.73
Mistral 7B 0.20 0.23 0.53 0.73 0.27 0.27 0.50 0.73
Gemma 7B (it) 0.25 0.26 0.58 0.73 0.25 0.26 0.49 0.73
Gemma 7B 0.41 0.37 0.50 0.27 0.19 0.24 0.49 0.73
Gemma 2B (it) 0.73 0.73 0.52 0.27 0.02 0.20 0.50 0.73
Gemma 2B 0.25 0.26 0.51 0.34 0.27 0.27 0.50 0.73
LR 0.02 0.19 0.61 0.74 0.02 0.19 0.61 0.74
XGBoost 0.00 0.16 0.74 0.76 0.00 0.16 0.74 0.76
@@
Table A2: Zero-shot LLM results on the ACSMobility benchmark task, together with supervised learning baselines fitted on 0.6M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.39 0.40 0.65 0.55 0.15 0.27 0.58 0.57
Mixtral 8x22B (it) 0.31 0.33 0.66 0.59 0.12 0.24 0.64 0.59
Mixtral 8x22B 0.20 0.28 0.63 0.44 0.30 0.34 0.57 0.58
Llama 3 70B (it) 0.15 0.24 0.70 0.60 0.12 0.24 0.64 0.53
Llama 3 70B 0.09 0.24 0.67 0.55 0.08 0.25 0.52 0.46
Mixtral 8x7B (it) 0.45 0.45 0.66 0.52 0.09 0.24 0.61 0.57
Mixtral 8x7B 0.28 0.32 0.60 0.44 0.07 0.25 0.57 0.58
Yi 34B (it) 0.35 0.36 0.65 0.56 0.06 0.25 0.50 0.44
Yi 34B 0.08 0.24 0.62 0.56 0.14 0.27 0.53 0.44
Llama 3 8B (it) 0.19 0.28 0.60 0.57 0.11 0.25 0.56 0.56
Llama 3 8B 0.08 0.25 0.53 0.56 0.12 0.26 0.48 0.44
Mistral 7B (it) 0.41 0.42 0.59 0.57 0.11 0.25 0.55 0.56
Mistral 7B 0.05 0.25 0.57 0.56 0.44 0.44 0.50 0.56
Gemma 7B (it) 0.42 0.43 0.53 0.56 0.10 0.26 0.49 0.44
Gemma 7B 0.04 0.24 0.61 0.58 0.03 0.25 0.52 0.55
Gemma 2B (it) 0.34 0.36 0.49 0.56 0.19 0.28 0.50 0.56
Gemma 2B 0.09 0.26 0.48 0.44 0.44 0.44 0.50 0.56
LR 0.04 0.24 0.58 0.56 0.04 0.24 0.58 0.56
XGBoost 0.02 0.19 0.77 0.70 0.02 0.19 0.77 0.70
@@
Table A3: Zero-shot LLM results on the ACSTravelTime benchmark task, together with supervised learning baselines fitted on 1.3M samples.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.33 0.34 0.71 0.60 0.10 0.20 0.68 0.73
Mixtral 8x22B (it) 0.24 0.25 0.70 0.72 0.04 0.18 0.71 0.75
Mixtral 8x22B 0.32 0.30 0.59 0.30 0.29 0.29 0.54 0.70
Llama 3 70B (it) 0.16 0.21 0.69 0.75 0.13 0.20 0.73 0.75
Llama 3 70B 0.18 0.22 0.67 0.63 0.12 0.21 0.64 0.53
Mixtral 8x7B (it) 0.20 0.23 0.70 0.74 0.06 0.19 0.69 0.74
Mixtral 8x7B 0.41 0.37 0.57 0.30 0.20 0.25 0.56 0.70
Yi 34B (it) 0.06 0.19 0.67 0.74 0.22 0.24 0.57 0.31
Yi 34B 0.04 0.21 0.59 0.70 0.09 0.20 0.67 0.64
Llama 3 8B (it) 0.11 0.21 0.59 0.71 0.17 0.22 0.64 0.68
Llama 3 8B 0.41 0.38 0.55 0.30 0.20 0.25 0.51 0.34
Mistral 7B (it) 0.30 0.30 0.61 0.70 0.07 0.20 0.67 0.65
Mistral 7B 0.29 0.30 0.45 0.30 0.30 0.30 0.50 0.70
Gemma 7B (it) 0.30 0.34 0.46 0.50 0.18 0.24 0.57 0.61
Gemma 7B 0.15 0.23 0.49 0.49 0.18 0.26 0.48 0.70
Gemma 2B (it) 0.70 0.70 0.54 0.30 0.24 0.29 0.42 0.42
Gemma 2B 0.26 0.28 0.54 0.30 0.30 0.30 0.50 0.70
LR 0.03 0.19 0.70 0.72 0.03 0.19 0.70 0.72
XGBoost 0.00 0.14 0.84 0.80 0.00 0.14 0.84 0.80
@@
Table A4: Zero-shot LLM results on the ACSPublicCoverage benchmark task, together with supervised learning baselines fitted on 1.0M samples.
(a) Multiple-choice prompting.
(b) Numeric prompting.
Figure A4: Calibration curves for the ‘GPT 4o mini’ model, across different race sub-populations on the ACSIncome task, computed using $10$ quantile-based bins. This model is only available through a web API, and no base model variant is available. Numeric prompting (Fig. 4(b)) leads to reduced differences in group-conditional calibration (improved group fairness), as well as large improvements in overall calibration.
Figure A5: Calibration curves across different race sub-populations on the ACSIncome task, computed using $10$ quantile-based score bins, with $95\%$ confidence intervals. Using multiple-choice prompting to generate risk scores (left plots) leads to sizeable calibration error for both base and instruction-tuned models. The models shown in this figure generally do not follow the same group bias trend seen in the Mixtral 8x22B and Yi 34B models (Figure 6).
A.3 Additional subgroup calibration results
@@
This section contains additional subgroup calibration plots on the ACSIncome task. Figure 6 shows the 2 worst offenders: Mixtral 8x22B and Yi 34B. Figures A4 and A5 show the remaining models. The same trend is visible but to a lesser extent: Instruction-tuned models show a bias towards lower scores for Black individuals, even after controlling for label prevalence.
@@
Figure A6 shows the difference between the signed calibration error (SCE) on different group pairings, on the ACSIncome task. Positive predictions ($Y=1$) correspond to the advantaged high-income class. Negative differences, $\Delta_{SCE}<0$, indicates advantaged scores for Black individuals, while positive differences, $\Delta_{SCE}>0$, indicate disadvantaged scores for Black individuals. Interestingly, while subgroup calibration curves appear similar for base models and disadvantage Black individuals for instruction-tuned models (see Figures 6A4, and A5), this is not entirely reflected on the SCE metric. Indeed, a clear-cut split is visible: base models benefit the score of Black individuals, and instruct models disadvantage the score of Black individuals. This finding could be partially explained by the fact that base models produce score distributions with low variance, and instruct models produce high-variance polarized outcomes. Specifically, two conclusions can be drawn from the score distributions produced by base models: (1) models under-estimate the score of high-income earners (which are disproportionately Asian), and (2) models over-estimate the score of low-income earners (which are disproportionately Black). This could lead base models to over-estimate the earnings of the Black population disproportionately to other groups. The opposite is true for instruction-tuned models: high-income earners see their score over-estimated, which benefits groups with a higher prevalence of high earners.
@@
Figure A6: Racial group bias in risk score calibration error. When comparing score bias of group $A$ with score bias of group $B$, $\Delta_{\mathrm{SCE}}=\mathrm{SCE}_{A}-\mathrm{SCE}_{B}$, positive values indicate an undue score advantage of group $A$, and negative values an undue score advantage of group $B$. Left: Difference between White and Black groups, $\mathrm{SCE}_{White}-\mathrm{SCE}_{Black}$. Right: Difference between Asian and Black groups, $\mathrm{SCE}_{Asian}-\mathrm{SCE}_{Black}$. Note that the Gemma models were omitted, as their instruct versions degenerate into strictly predicting the same outcome for all samples. Consequently, the two instruct Gemma models are the only exceptions to the trend shown in these plots.
A.4 Varying uncertainty
Figure A7: Shift of score distribution with increasing evidence for the Mixtral 8x7B model (which achieved the best Brier score), using multiple-choice Q&A on the ACSIncome task. Features are described in Table A5. Score distribution gets more discriminative as more evidence is added, successfully increasing scores’ predictive signal (AUC). The true label prevalence is $\mathbb{P}[Y=1]=0.37$.